OpenStudy (anonymous):

A sail is in the form of a right triangle that is three times as high as it is wide. The sail is made from 24 square meters of material. WHAT IS ITS HEIGHT??? Remember...a=1/2bh

OpenStudy (anonymous):

Use pythogorean theorem i guess.

OpenStudy (anonymous):

dont know what that means lol

OpenStudy (anonymous):

area, A = 24 = (1/2)bh and the problem tells you that height is 3 times width (which is base), so h = 3w = 3b

OpenStudy (anonymous):

? h=3w=3b?

OpenStudy (anonymous):

??????? that doesnt make sense to me can you help me more with how to solve that again ugh

OpenStudy (anonymous):

Area = 0.5 * Base * Height The question tells you that the height is 3 times the base. substitute what you know into the equation and you should only have one variable left. Solve for that variable.

OpenStudy (anonymous):

 Area=1/2 (x) (3x) =3/2 x^2 =24 --->x^2=16...x=4...height =3x...

OpenStudy (anonymous):

@elizabeth.a.fairweather do you get it

OpenStudy (anonymous):

@elizabeth.a.fairweather height =3x =12

OpenStudy (anonymous):

yes that does make sense thank you